How many photons are produced in a laser pulse of .125 J at 523 nm
Answer #1

Energy= hc/lambda

H= 6.63*10^-34 J/s

C= 3*10^8 m/s

Wavelength=523nm = 523*10^-9 m

E=( 6.63*10^-34)3*10^8/523*10^-9 =3.8*10^-19 J

So 0.125 J * ( 1photon/3.8*10^-19)

= 3.28*10^17 photons